rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#8179 — FS#12050 — Lille Optic Loop

Attached to Project— ADSL / SDSL
Maintenance
Backend / Core
CLOSED
100%
Maintenance on the fiber optics is being performed by the operator in Lille, no impact for clients.
Date:  Tuesday, 25 November 2014, 12:59PM
Reason for closing:  Done
Comment by OVH - Tuesday, 18 November 2014, 10:05AM

The optic loop is secure again.
There will be ongoing maintenance tonight from midnight to 6am, still no impact to customers.